Calculate the molarity for each solution.

87.2 g of Na2SO4 in enough water to make 500. mL of solution

61.8 g of NH3 in enough water to make 7.00 L of solution

100. mL of ethanol (C2H5OH) in 500. mL of solution (The density of ethanol is 0.789 g/mL.)
Answer #1

1) Molarity(M) = (w/Mwt)*(1000/v)

w = weight of solute(Na2SO4) = 87.2 g

Mwt = molarmass of Na2SO4 = 142 g/mol

V = volume of solution = 500 ml

Molarity (M) = (87.2/142)*(1000/500)

              = 1.23 M

2) Molarity(M) = (w/Mwt)*(1/v)

w = weight of solute(NH3) = 61.8 g

Mwt = molarmass of nh3 = 17 g/mol

V = volume of solution = 7 l

Molarity (M) = (61.8/17)*(1/7) = 0.52 M

3) Molarity(M) = (w/Mwt)*(1/v)

w = weight of solute(C2H5OH) = 100*0.789 = 78.9 g

Mwt = molarmass of C2H5OH = 46 g/mol

V = volume of solution = 0.5 l

Molarity (M) = (78.9/46)*(1/0.5) = 3.43 M